OpenStudy (anonymous):

(4-m^2)/(3m-6) "m" not equal 2 simplify

OpenStudy (zzr0ck3r):

"m" not equal 2 simplify?

OpenStudy (campbell_st):

just factorise the numerator and denominator... then cancel the common factor Numerator (4- m^2) = -(m^2 - 4) = -(m -2)(m+2) denominator (3m - 6) = 3(m - 2) cancel the common factor

OpenStudy (zzr0ck3r):

(4-m^2) = (2^2-m^2) = (2-m)(2+m) this is form: (a^2-b^2) = (a-b)(a+b) and (3m-6) = 3(m-2) put it all together and we have (2-m)(2+m)/(3(m-2)) factor a negative out of the top -(m-2)(2+m)/(3(m-2)) = -(2+m)/3

OpenStudy (zzr0ck3r):

as long as m is not equal to 2...I see what you were saying.

OpenStudy (anonymous):

thanks